E. Sodium-24 (t - 15 h) is used to study blood circulation. If a patient is injected with an aqueous solution of
1
Expert's answer
2021-09-30T03:05:04-0400

Na-24 half life = 15h


We know that for first order reaction


"t \\frac{1}{\n\n2}\n\n\n\n\u200b\t\n\n = \n\n\n\n\\frac{0.693}{K}"


"Therefore, K = \\frac{0.693}{t\\frac{1}{2}} \n\n\u200b"


"K= \frac{0.693}{15h}= 0.0462h^{-1}"


At starting, (Activity)0 = 2.5x109d/s




At 4days, 4×24hr= 96hr


96×2.5×109


= 2.4×1011 d/s